OpenStudy (anonymous):

a 65cm tall graduated cylinder is filled with 30cm of glycerin (density 1260 kg/m^3) and 35 cm of water. find the pressure difference between the top and bottom of the cylinder.

OpenStudy (fretje):

P = rho.g.h (Pressure = density.g_gravitational acceleration.h_height) here we can add the partial pressures P = P1 + P2 = rho1.g.h1 + rho2.g.h2 = 30cm.1m/100cm.1260kg/m^3 . 9,81m/s^2 + 35cm.1m/100cm . 1000kg/m^3 . 9,81m/s^2 = 7141,68 Pa ?